UT 20130326 Technical Observing

Observer: DMiller, JHill (Tucson), JGuerra, JChristou (LBTO)
Telescope Operator: GBechetti (LBTO)
Additional AO Support: DMiller (LBTO)
Telescope Support: JLittle (LBTO)
Instrument Support: ESolheid (LBTO)

Summary

With both big "W" units broken, tonight has been reprogrammed from Adaptive Optics to Active Optics and Binocular activities.

We took fast (by little "w" standards) wfs data for Juan Carlos' experiment to measure the outer scale by comparing the two wavefronts (200 ms exposures at ~1 sec intervals).

We worked the last major bugs out of Doug's binocular preset scripts and tested them on-sky.

Then we started collection binocular pointing and collimation data with. This also serves as a check of the pointing models with LM2 and other ARGOS platforms installed.

The seeing got really nice, so we switched from pointing data to field aberration data. After two batches of field aberration data, we spent time trying to understand why the Issue 4545 cross-talk is different on the two sides. Guider images on SX were briefly down to 0.25 arcsec, and the running average to 0.30. We never got DX better than 0.4.

At the end the DX secondary failed. We were able to collect some more Zernike Gain Data on the SX side.

The wfs FWHM's seem to bottom out at a value of about 0.55 arcsec FWHM in good seeing (seeing on LFBG guider was 0.3).

Preparations

Doug drove down to Tucson.

John changed sextractor BACK_SIZE from 64 to 32 (for guiding). We are using BACK_SIZE=10 for wfs - see Issue 4490).

Details

Monocular AGw1

Binocular Presets with IDL

Doug's scripts are working OK - we can send binocular syncronous and async presets, but they need some work to be user friendly (even for Doug).

JC's Dual wfs experiment

02:55 Doing this for ~30 minutes while Doug prepares for binocular pointing and collimation. Seeing is ~1.1 arcsec. The idea is to read the little "w" sensors very fast and use the spot motions to measure the outer scale.

WT10_335 (too bright)

03:01 WT10_335 gs=1 (mag 8.6)

Manually set exposure times after the preset: wfs 1000 ms guider=250 ms
These data still being taken with the GCS guiding and wfsing loops as we collimate.

03:14 Clear active optics on Right M1

03:17 Clear active optics on Right M1

03:33 StopWFsing using buttons on GCS GUIs, then in two windows:

readWFSCam left -e 200 500 AND readWFSCCam right -e 200 500 more or less simultaneously (limited by Polycom delay)

Data in /home/lbto/data/FastWFS/

200 ms full frame images are at approx 1 sec intervals.

03:41 Right primary mirror panics around image 184 (took 187). Why?

Left data continued until about image 250 (took 264)

03:47 resend ACTIVE presets on both sides

03:50 GCS recovered after a "failed to receive image"

03:54 Right M1 is at force limit. Clear active optics on Right M1

03:56 Active Optics has driven in a bunch on Z11/Z22 on right.

04:04 Start another set of 500 fastWFS data. Right starts at 000188, Left starts at Seeing on guider is 1.2 on left 1.3 on right (slightly worse collimation and more dome seeing).

04:26 Right side finishes the 500 images first (by 4 minutes). Those old shuttle boxes on AGW2 are still holding their own!

Binocular Collimation and Pointing Data

04:38 Point to BS9142 star for collimation at normal speed. Right guide star stars at 0.7 FWHM without any collimation. On-axis collimation seems happy here on both sides. Seeing varying from 0.52 to 0.72 arcsec FWHM on guider.

04:57 First try at binocular pointing script, but Doug isn't sending the SYNCPRESET.

05:14 SYNC preset succeeds from ptmodel_binoc_collect.pro.

05:15 Turn on pointing logs

05:19 Right shell rips during the slew to the second star. 1 minute delay before IIF knows this and cancels the preset.

05:38 Corrected pointing on both sides.

05:42 editted PCSInstrument.conf to make the non-LBTI gregorian LEAKRATES 2.5 (most of them were already there - LBTI uses 5.0)

Thin clouds are moving in from the northwest.

06:06 Seeing is 0.5 arcsec at EL=43.5 degrees, so we will interrupt this pointing model data for some low-elevation field aberration data.

Field Aberration Data on BS9155

06:15 SYNC Preset to BS9155.

06:30 field_collect with async preset on left side. Seeing is 0.6 arcsec on guider at EL=33 through thinnish clouds.

06:33 field_collect with async preset on right side, so now we are collecting field data at twice the rate!

We can see that DX has some field aberration. That's no surprise as M2 RY is at 530 arcsec.

07:13 Seeing continues to be good with 0.6 arcsec on the guider at EL=41 deg.

07:33 This field data collection is complete.

07:34 Stop and restart OSS to fix up telemetry (per Stephen).

07:35 The really good seeing seems to have ended.

Field Aberration Data on BS9152

07:50 Preset to higher elevation for another set of field data on both sides. BS9152

Adjust pointing on Right

07:57 field_collect started on left side. Seeing on guider is 0.5, but less impressive as we are at EL=67.

08:xx field_collect started on right side.

09:02 This field data collection is complete.

DX Hotspot adjustment

Previous 310, 239

09:06 300, 239 (Z11 went down from ~350 to ~-50)

09:11 310, 239 (Does Z11 go back up?)

09:17 Clear Active Optics on DX M1 and stop reporting Zernikesfor ~ 1 min

09:18 298, 239

09:20 297, 243

09:22 287, 245

09:28 295,245

09:32 295,243 (where John's eye thinks the hotspot should be)

Saved that hotspot in LUCI_IRTC_R.cfg

Adjusting reconstructor masks on DX

DX still has this high order wf stuff going on. I think that's why it has 0.45 on the guider while SX has 0.35.

09:44 change mask13.msk to std_mask.msk in LUCI_IRTC_R.cfg

09:47 mask9x9.msk

Geno focussed the DIMM and made it better.

09:50 back to the normal mask13.msk since those only work for CoW

Down to 0.23 on SX guider.

09:58 Remove -200 Z12

10:00 change from 9x9 "20121213" to 11x11 reconstructor "20120816"

Manual Global Offsets of Z11 Z22

Block DX Zernikes

10:04 Manual adjustment of Z22 DX global offset down to -260 in steps

10:13 DX Zernikes back on

Block SX Zernikes

10:14 -100 Z22 SX global

10:16 -200 Z22 SX global

10:17:30 0 Z22 SX global

10:18 -500 Z11 SX global produces Z4, Z11, Z22

10:22 Enable SX Zernikes, Block DX Zernikes

10:22 -500 Z10 on right

10:24 0 Z10

Enable RIGHT

10:30 Back to 9x9 on Right 20121213

10:32 -500 Z11 DX global

10:34 all DX globals to zero

Zernike Gain Data

10:39 Preset to BS9162 - Shell RIP'd during preset

10:46 start collecting zernike data on the left

10:52 DX shell RIP'd again. Doug is cycling the power. But the problems persist........... Data collection continues on the left side.

11:05 start multi_step_focus since DX is cancelled.

Some consultation with Runa gets DX AdSec running again.

11:22 Clear Active Optics on Both Primaries

11:27 start collecting shortened zernike data on the right

11:28 DX shell RIP'd again.

11:44 DX Shell RIP'd again. (He removed it from the position list, but not the current list.)

12:16 left zernikes data is finished

Still problems with DX M2.

Close

12:25 close for impending dawn. Seeing on SX guider is 0.33 FWHM.

-- JohnHill - 26 Mar 2013
Topic revision: r5 - 17 May 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback